three charges of 2mC 
,-4mC and -4mC are kept at vertices of euilateral traingle of side 10 cm
. a charge q is kept on the mid point of -4mC  and -4mC charge.calculate the value of q such that 2mC charge remains in equilbrium.
 

Grade:12

1 Answers

Khimraj
5 years ago
net force on 2mC = 0
so
2*(K*2mC*(-4mC)/0.12)cos30 + K*2mC*q/(3*0.12/4) = 0
-8mC*cos30 = -4q/3
q = 6mC*cos30 = 3mC
Hope it clears............................